To prevent the presence of air, noble gases are placed over highly reactive chemicals to act as inert “blanketing” gases. A chemical engineer places a mixture of noble gases consisting of 5.50 g of He, 15.0 g of Ne, and 35.0 g of Kr in a piston-cylinder assembly at STP. Calculate the partial pressure of each gas


1
Expert's answer
2021-12-14T13:00:02-0500

Dalton's Law of partial pressure

Ptotal = P(He) + P(Ne) + P(Kr)

P(He) = X(He)*Ptotal, P(Ne) = X(Ne)*Ptotal, P(Kr) = X(Kr)*Ptotal

X(He)= n(He)/ntotal, X(Ne)= n(Ne)/ntotal , X(Kr)= n(Kr)/ntotal

ntotal = n(He) + n(Ne) + n(Kr)

Calculations:

n= m/M

n(He) = 5.50 g/ 4.003 g/mole = 1.374 mol

n(Ne) = 15.0 g / 20.18g/mole = 0.7433 mol

n(Kr) = 35.0 g/83.80 g/mole = 0.4177 mol

ntotal = 1.374 mol + 0.7433 mol + 0.4177 mol = 2.535 mol

P(He) = X(He)*Ptotal = (1.374 mol/2.535mol)(1 atm) = 0.5420 atm

P(Ne) = X(Ne)*Ptotal = (0.7433 mol/ 2.535 mol)(1 atm) = 0.2932 atm

P(Kr) = X(Kr)*Ptotal = (0.4177 mol/2.535 mol)*(1 atm) = 0.1648 atm

Answer: P(He) = 0.5420 atm

P(Ne) = 0.2932 atm

P(Kr) = 0.1648 atm